EMISSION SPECTRA OF H_{2}^{17}O AND H_{2}^{18}O FROM 320 TO 2500 CM^{−1}

\underline{SEMEN MIKHAILENKO}, Laboratory of Theoretical Spectroscopy, Institute of Atmospheric Optics, 1, Akademicheskii Av., 634055 Tomsk, Russia; GEORG Ch. MELLAU, Justus-Liebig-Universität Giessen, Physicalisch-Chemisches-Institut, Heinrich-Buff-Ring, 58, 35392 Giessen, Germany; and VLADIMIR G. TYUTEREV, Groupe de Spectroscopie Moléculaire et Atmosphérique, UMR CNRS 6089, UFR Sciences, BP 1039, 51687 Reims Cedex 2, France.

Emission spectra of water enriched by oxygen-17 and by oxygen-18 have been recorded in the 322-850, 1136-1790, and 1654-2500 cm^{−1} spectral regions with pressures 1-18 mbar and temperature up to 1950 K using Bruker IFS 120 HR spectrometer. The measurements were performed in an alumina cell with an effective length of hot gas of about 50 cm. The relative concentrations of the H_{2}^{16}O, H_{2}^{17}O, and H_{2}^{18}O species were from 75/15/10 up to 50/30/20 percents respectively for different spectra. A lineshape analysis using the SpectrumFit program^{a} was done for a precise determination of the line positions. Totally the line parameters of more then 12000 water lines were derived. Main part of observed lines assigned to transitions of the (000), (010), (020), (100), (001), (030), (110), and (011) states of mentioned above water isotopomers. New results for the line positions of different ro-vibrational bands and energy levels of the H_{2}^{17}O and H_{2}^{18}O molecules are presented. A comparison of the observed line positions with previous experimental and theoretical studies is discussed.

ahttp://www.spectrumfit/symath.com